OpenStudy (anonymous):

x=2t^2(e^2T)-2te^2t+e^2t differentiate and simplify

OpenStudy (john_es):

You should use the chain law of differentation and the rule for the sum. For example, suppose you have, \[x(t)=te^{2t}+t^2\]Then, \[(x(t))'=(te^{2t}+t^2)'=\\ =(te^{2t})'+(t^2)'=(t)'e^{2t}+t(e^{2t})'+(t^2)'=\\ =e^{2t}+2te^{2t} +2t\] Try with your problem.

OpenStudy (anonymous):

thanks